An object moves with constant acceleration 4.70 m/s2 and over a time interval reaches a final velocity of 13.6 m/s.

(a) If its original velocity is 6.80 m/s, what is its displacement during the time interval?


(b) What is the distance it travels during this interval?


(c) If its initial velocity is −6.80 m/s, what is its displacement during this interval?


(d) What is the total distance it travels during the interval in part (c)?
Answer #1

Given :- a = 4.7 m/s?, final velocity V=136 mis a) initial velocity = 680 mis displacement ? 6.domis 13.6 m/s a = 4.7 m/s² tfor AB 1154, -6.80ms 9:491 mis? t> 0-4 , V-0 = 42 tnt, , SES; tj = 0 - (-68) od 4465 Lp tra luss Si = lt + at? = litit. I ator we com also write total displacement as rector sum, with direction (sign) 8 = sits = 1965 – 4892 13= 14.73m) long A to d)